a rectangle is 6 feet long and 4 feet wide. if each dimension is increased by the same number of feet, a new rectangle is formed whose area is 39 square feet more then the area of the original rectangele. by how many feet was each dimension increased?
Answer by checkley77(12844)

AREA=6*4=24 FT^2
39+24=(6+X)(4+X)
63=24+10X+X^2
X62+10X+24-63
X^2+10X-39=0
(X+13)(X-3)=0
X-3=0
X=3 FT. ANS. FOR THE INCREASE IN LENGTH & WIDTH.
pROOF:
63=(6+3)(4+3)
63=9*7
63=63